Substitution as suggested by the equation

sinxsinydx + cosxcosydy = 0


1
Expert's answer
2022-03-28T05:37:44-0400

"\sin x \sin y dx+\cos x \cos y dy=0"

"\frac{\sin x }{\cos x} dx=-\frac{\cos y}{ \sin y} dy"

"-\frac{d\cos x }{\cos x}=-\frac{d\sin y}{ \sin y}"

"\ln|\cos x|=\ln|\sin y|+\ln C"

"\cos x=C\sin y"

Answer: "\cos x=C\sin y".